# bibfiles/main.bib

### v1

@@ Abstract
Human Activity Recognition (HAR) with wearable sensors is challenged by limited interpretability, which significantly impacts cross-dataset generalization. To address this challenge, we propose Motion-Primitive Transformer (MoPFormer), a novel self-supervised framework that enhances interpretability by tokenizing inertial measurement unit signals into semantically meaningful motion primitives and leverages a Transformer architecture to learn rich temporal representations. MoPFormer comprises two-stages. first stage is to partition multi-channel sensor streams into short segments and quantizing them into discrete “motion primitive” codewords, while the second stage enriches those tokenized sequences through a context-aware embedding module and then processes them with a Transformer encoder. The proposed MoPFormer can be pre-trained using a masked motion-modeling objective that reconstructs missing primitives, enabling it to develop robust representations across diverse sensor configurations. Experiments on six HAR benchmarks demonstrate that MoPFormer not only outperforms state-of-the-art methods but also successfully generalizes across multiple datasets. Most importantly, the learned motion primitives significantly enhance both interpretability and cross-dataset performance by capturing fundamental movement patterns that remain consistent across similar activities regardless of dataset origin.
@@ 1 Introduction
Human Activity Recognition (HAR) has a wide range of applications and can be achieved through various methods [sedaghati2025application, arshad2022human]. While vision-based and environmental sensor-based methods have been extensively explored [hussain2019different, al2020review], systems leveraging Inertial Measurement Units (IMUs) offer distinct advantages, including compact size, low power consumption, and minimal computational overhead for data acquisition and preprocessing. IMUs also inherently preserve privacy compared to visual alternatives [chen2025comodo, padmanabha2025egocharm] and are robust to environmental factors like lighting variations and occlusions [kamboj2024survey]. Such characteristics make IMUs a highly suitable sensing modality for capturing human motion across diverse real-world scenarios.
@@
Despite these strengths, IMU-based HAR confronts a significant hurdle: interpretability. IMU data, comprising multi-channel time series of linear acceleration and angular velocity, is inherently less intuitive for humans to interpret than video data, making it hard to understand what the model has learned [geissler2024strategies]. Moreover, many human activities exhibit hierarchical structures, composed of more fundamental activities or movements. However, there is a common absence of fine-grained labels for the sub-activities that constitute these complex, hierarchical activities. For example, an activity like “washing hands” can be deconstructed into sub-activities such as “turning on the tap”, “applying soap”, scrubbing hands”, and “rinsing hands”. A HAR model might learn to identify “washing hands” by detecting the “turning on the tap” sub-activity, especially if it is a prominent signal. However, if the model encounters a different activity, such as “filling a water bottle”, which also begins with “turning on the tap”, it may struggle for HAR models to differentiate between the two activities. This is because the model might rely on the common initial sub-activity rather than grasping the complete, semantically meaningful sequence (e.g., the subsequent application of soap and scrubbing unique to handwashing) that defines the core of the activity. Each of these sub-activities might, in turn, be composed of even finer-grained movements. The black-box nature of contemporary HAR models often obscures whether these models are learning the underlying activity structure or merely exploiting superficial data correlations, thereby exacerbating the interpretability challenge [lipton2018mythos, arabzadeh2025cnn].
@@
Another challenge in sensor-based HAR is data heterogeneity. Variations in sensor types, sampling rates, on-body placement (e.g., sensors worn on the wrist or the waist), subject characteristics (e.g., age, height, fitness level of the individual), and environmental conditions can introduce significant distributional shifts [panahandeh2013continuous]. This heterogeneity impedes model generalization across different contexts. Consequently, developing strategies to address data heterogeneity is crucial for building robust HAR systems.
@@
To address these challenges, we introduce Motion-Primitive TransFormer (MoPFormer), a novel pre-training architecture for IMU-based HAR that significantly enhances model interpretability. Inspired by advancements in language modeling [grattafiori2024llama], MoPFormer conceptualizes IMU motion sequences as a series of “Motion Primitives” that serve as semantically meaningful, discrete building blocks of human activity. As illustrated in Fig. 1, our approach first divides raw IMU sequences into short segments, analogous to “words” in natural language, creating an interpretable vocabulary of fundamental movements. Each segment integrates data from various sensor channels at concurrent time points, forming feature vectors that are then arranged sequentially into a two-dimensional feature matrix. This construction not only mirrors the arrangement of words in a sentence but also enables transparent analysis of which primitive movements contribute to activity recognition decisions. To further enrich semantic understanding, we embed metadata for each sensor using a context-aware embedding module. The resulting representation allows for direct examination of motion primitive similarities, frequencies, and their distribution across different activities, providing unprecedented insights into model behavior. This comprehensive representation is subsequently processed by a Transformer backbone [vaswani2017attention], enabling MoPFormer to effectively handle heterogeneous channel configurations and learn robust, interpretable features for diverse downstream tasks.
@@
  • We propose MoPFormer, an effective pre-training framework that achieves state-of-the-art performance in our experiments.
@@
  • We extract motion primitives and provide a detailed interpretability analysis of these primitives.
@@ 2 Related Work
Traditional human activity recognition approaches train model parameters on different datasets separately, including statistical feature extraction methods and deep learning approaches. Deep learning methods include CNN-based architectures such as CALANet [park2024calanet], COA-HAR [xu2022shallow], MA-CNN [3161174], and SenseHAR [3360032], RNN-based approaches like DeepConvLSTM [ordonez2016deep], and attention-based models such as THAT [li2021two] and PA-HAR [pmlr-v189-xu23a]. Recent research has proposed general-purpose time series models applicable to various tasks, including classification, such as TimesNet [wu2022timesnet], TSLANet [10.5555/3692070.3692564], and FITS [xu2023fits]. However, these methods are typically trained and tested on splits from the original datasets and are limited in their ability to achieve cross-dataset generalization.
@@
Self-supervised human activity recognition methods perform representation learning through various approaches. Reconstruction-based methods such as TST [zerveas2021transformer] and TARNet [chowdhury2022tarnet] focus on rebuilding input signals. Contrastive learning methods such as TS2Vec [yue2022ts2vec], CL-HAR [3539134], DDLearn [radford2021learning], TS-TCC [eldele2021time], FOCAL [liu2023focal], and ModCL [10475144] execute discriminative representation learning tasks. Other self-supervised objectives like BioBankSSL [zerveas2021transformer, doherty2017large] and Step2Heart [spathis2021self] also contribute to learning time series representations. These methods typically learn feature representations on specific datasets and fine-tune classifier heads on the target dataset, but they can only perform on datasets with similar structures and still require labeled data from the target dataset.
@@
Research on sensor-based HAR for composite activity recognition is relatively limited [chen2021deep]. For instance, Chen et al. decomposed composite activities into multiple simple activities [chen2021deep], where each simple activity is represented by a sequence of sensor signal segments. These segments are first fed into a CNN to extract representations for identifying simple activities. Concurrently, the CNN-extracted features from all segments are passed to an LSTM network to achieve high-level semantic activity classification. Similarly, prior work [ijcai2018p461] inferred composite activities using estimated activity sequences, where temporal correlations of simple activities were extracted for composite activity classification. Conversely, predicted composite activities were used to aid the derivation of simple activity sequences for the next time step, with predictions for both simple and composite activity sequences mutually updating during inference. While these studies present viable approaches for composite activity research, datasets with comprehensive, multi-level hierarchical labels for composite activities are scarce, and existing hierarchical annotations are often incomplete.
@@
The concept of motion primitives has gained traction in robotics research [sanzari2019discovery, 10.5555/3692070.3692298]. These primitives are fundamental movement patterns that can be sequenced and combined to generate or understand a diverse repertoire of complex human activities. For instance, Calvo et al. utilized eight such primitive movements, including “static” and partial “squat”, to train Hidden Markov Models for activity classification. Similarly, Sanzari et al. developed a framework for automatically discovering human motion primitives from 3D pose data by optimizing “motion flux” and then clustering the results [sanzari2019discovery]. However, the application of motion primitives to enhance the interpretability of IMU-based HAR has seen limited progress. This limitation stems largely from the inherently non-intuitive nature of IMU sensor data, which makes it challenging to define and interpret primitives directly from these signals.
@@
Much of the research on interpretability in HAR has focused on attention mechanisms. The principle behind deep attention models is to weight input components, with the assumption that components assigned higher weights are more relevant to the recognition task and exert greater influence on the model’s decisions [serrano2019attention]. For example, Shen et al. [shen2018sam] designed a segment-level attention method to determine which time periods contain more information; combined with a gated CNN, this segment-level attention can extract temporal dependencies. Zeng et al. [zeng2018understanding] developed attention mechanisms from two perspectives: they first proposed sensor attention on the input to extract salient sensory patterns, and then applied temporal attention to LSTMs to filter out inactive data segments. Ma et al. [ma2019attnsense] employed spatial and temporal attention mechanisms, extracting spatial dependencies by fusing patterns with self-attention. However, their interpretability often remains at a superficial level, indicating what the model focuses on, rather than why or how these features contribute to the recognition of complex, hierarchically structured activities.
@@ 3 Methodology
The architecture of MoPFormer, as illustrated in Fig. 2, comprises four modules: the motion primitive module, the context-aware embedding module, and two task heads. The motion primitive module first transforms multi-dimensional time-series data windows into index sequences from a fixed vocabulary. To further extract patterns, the context-aware embedding module extracts semantic information from contextual relationships and embeds high-dimensional vector sequences for the Transformer backbone. The MAE head, as a pre-training task, strengthens the context-aware embedding module’s ability to understand motion primitives, while the CLS head enables the Transformer backbone to learn effective features for classification. Fig. 2 illustrates the detailed architecture of each key module in our framework.
Figure 2: Detailed illustration of key modules in our motion-centric framework. (a) Motion Primitive Module: raw data windows from multiple sensor channels are partitioned into motion primitives and processed through instance normalization to generate VQ indices and statistical features. (b) Context-Aware Embedding Module: special tokens are inserted alongside masked motion embedding tokens ($M$) to form the complete input representation, which combines motion primitive embeddings, positional encodings, and statistical feature embeddings. (c) Task Heads: The diagram shows transformed features $X^{*}$ representing the corresponding position vectors after Transformer Encoder processing. The MAE head utilizes $M^{*}$ positions for masked token prediction during pretraining, while the trainable CLS head operates exclusively on the transformed [CLS] token representation for downstream task fine-tuning.
@@ 3.1 Motion Primitive Module
The Motion Primitive Module transforms raw data windows into discrete motion primitives, serving as the foundational, interpretable units for our motion-centric representation. For each data window of shape $T\times C$, where $T$ is the number of time steps and $C$ is the number of channels (e.g., accelerometer, gyroscope), we first partition the sequence into non-overlapping segments across each channel independently. Each channel is divided into segments of length $L$, resulting in $S=\lfloor T/L\rfloor$ segments per channel and a total of $S\times C$ segments per window.
Each segment undergoes a feature extraction process where we compute both low-level temporal patterns and statistical characteristics. For each segment $s_{i}\in\mathbb{R}^{L}$ (a single-channel segment), we apply instance normalization in Eq. (1) to standardize the input:
@@
The normalized segments are then encoded into discrete motion primitives using a Vector Quantization (VQ) [van2017neural] approach. We maintain a learnable codebook $\mathcal{Z}=\{z_{1},z_{2},...,z_{K}\}$, which contains $K$ prototype vectors (or “motion primitives”), each of dimension $L$. As illustrated in Fig. 2(a), for each normalized segment $\hat{s}_{i}$, we compute its VQ index through Eq. (2) by finding the nearest prototype in the codebook:
@@ 4.1 Experimental Setup
We conducted extensive evaluations using six publicly available benchmark datasets: PAMAP2 [pamap2_physical_activity_monitoring_231, Reiss2012IntroducingAN], DSADS [daily_and_sports_activities_256, Altun2010ComparativeSO], MHealth [mhealth_319, banos2015design], Realworld [7456521], UCI-HAR [human_activity_recognition_using_smartphones_240, Anguita2013APD], and USC-HAD [mi12:ubicomp-sagaware]. To ensure fair comparison with existing self-supervised methods, we pre-trained our model on five of these datasets while using the remaining dataset for evaluation, maintaining consistent experimental settings with other self-supervised approaches. All datasets are resampled to 100 Hz and segmented with a 500-sample window. To eliminate information leakage from overlapping windows, we adopt the setting by using strides equal to the window size for all labeled training segments. This strict non-overlapping approach may reduce absolute performance compared with more permissive settings, but ensures a more rigorous evaluation. We use accuracy and macro-F1 as evaluation metrics to assess model performance.
@@ 4.2 Results
BYOL [grill2020bootstrap]+perm_jit Acc 80.88 91.30 88.94 85.18 82.66 74.04 83.83
@@
BYOL [grill2020bootstrap]+Mixup Acc 82.38 93.87 90.06 88.80 85.31 75.59 86.00
@@
We compared MoPFormer with six different baseline approaches. Following the comprehensive evaluation in [3539134], we selected the BYOL framework [grill2020bootstrap], which demonstrated the best average performance among various contrastive learning frameworks. For time-domain and frequency-domain augmentations, we used perm_jit and lfc, respectively, which showed superior performance in their domains. As noted in [10177231], Mixup provides distinctive benefits compared to other temporal augmentation methods, so we included it as a baseline. ModCL [10475144] represents a state-of-the-art contrastive learning method that leverages both intra-modal and inter-modal consistency in wearable sensor data. We included it as a contemporary contrastive learning approach. For all contrastive learning methods, we maintained a consistent 0.2:0.8 train-test split ratio.
Additionally, we compared against recent supervised approaches, including TSLANet [10.5555/3692070.3692564] (ICML 2024), a widely recognized foundational model in the time-series domain, and CALANet [NEURIPS2024_8078e76f] (NeurIPS 2024), which was specifically designed for HAR tasks [BergmeirNeurIPS2024talk]. Detailed parameter configurations for all baseline models are provided in Appendix A.

### v2

# MoPFormer: Motion-Primitive Transformer for Wearable-Sensor Activity Recognition

@@ Abstract
Human Activity Recognition (HAR) with wearable sensors is challenged by limited interpretability, which significantly impacts cross-dataset generalization. To address this challenge, we propose Motion-Primitive Transformer (MoPFormer), a novel self-supervised framework that enhances interpretability by tokenizing inertial measurement unit signals into semantically meaningful motion primitives and leverages a Transformer architecture to learn rich temporal representations. MoPFormer comprises two stages. The first stage is to partition multi-channel sensor streams into short segments and quantize them into discrete “motion primitive” codewords, while the second stage enriches those tokenized sequences through a context-aware embedding module and then processes them with a Transformer encoder. The proposed MoPFormer can be pre-trained using a masked motion-modeling objective that reconstructs missing primitives, enabling it to develop robust representations across diverse sensor configurations. Experiments on six HAR benchmarks demonstrate that MoPFormer not only outperforms state-of-the-art methods but also successfully generalizes across multiple datasets. More importantly, the learned motion primitives significantly enhance both interpretability and cross-dataset performance by capturing fundamental movement patterns that remain consistent across similar activities, regardless of dataset origin.
@@ 1 Introduction
Human Activity Recognition (HAR) has a wide range of applications and can be achieved through various methods [42, 4]. While vision-based and environmental sensor-based methods have been extensively explored [23, 18], systems leveraging Inertial Measurement Units (IMUs) offer distinct advantages, including compact size, low power consumption, and minimal computational overhead for data acquisition and preprocessing. IMUs also inherently preserve privacy compared to visual alternatives [10, 31] and are robust to environmental factors like lighting variations and occlusions [25]. Such characteristics make IMUs a highly suitable sensing modality for capturing human motion across diverse real-world scenarios.
@@
Despite these strengths, IMU-based HAR confronts a significant hurdle: interpretability. IMU data, comprising multi-channel time series of linear acceleration and angular velocity, is inherently less intuitive for human to interpret than video data, making it hard to understand what the model has learned [19]. Moreover, many human activities exhibit hierarchical structures, composed of more fundamental activities or movements. However, there is a common absence of fine-grained labels for the sub-activities that constitute these complex, hierarchical activities. For example, an activity like “washing hands” can be deconstructed into sub-activities such as “turning on the tap”, “applying soap”, “scrubbing hands”, and “rinsing hands”. An HAR model might learn to identify “washing hands” by detecting the “turning on the tap” sub-activity, especially if it is a prominent signal. However, if the model encounters a different activity, such as “filling a water bottle”, which also begins with “turning on the tap”, it may struggle for HAR models to differentiate between the two activities. This is because the model might rely on the common initial sub-activity rather than grasping the complete, semantically meaningful sequence (e.g., the subsequent application of soap and scrubbing unique to handwashing) that defines the core of the activity. Each of these sub-activities might, in turn, be composed of even finer-grained movements. The black-box nature of contemporary HAR models often obscures whether these models are learning the underlying activity structure or merely exploiting superficial data correlations, thereby exacerbating the interpretability challenge [27, 3].
@@
Another challenge in sensor-based HAR is data heterogeneity. Variations in sensor types, sampling rates, on-body placement (e.g., sensors worn on the wrist or the waist), subject characteristics (e.g., age, height, fitness level of the individual), and environmental conditions can introduce significant distributional shifts [32]. This heterogeneity impedes model generalization across different contexts. Consequently, developing strategies to address data heterogeneity is crucial for building robust HAR systems.
@@
To address these challenges, we introduce Motion-Primitive TransFormer (MoPFormer), a novel pre-training architecture for IMU-based HAR that significantly enhances model interpretability. Inspired by advancements in language modeling [20], MoPFormer conceptualizes IMU motion sequences as a series of “Motion Primitives” that serve as semantically meaningful, discrete building blocks of human activity. As illustrated in Fig. 1, MoPFormer first divides raw IMU sequences into short segments, analogous to “words” in natural language, creating an interpretable vocabulary of fundamental movements. Each segment integrates data from various sensor channels at concurrent time points, forming feature vectors that are then arranged sequentially into a two-dimensional feature matrix. This construction not only mirrors the arrangement of words in a sentence but also enables transparent analysis that which motion primitives contribute to activity recognition decisions. To further enrich semantic understanding, we embed metadata of each sensor using a context-aware embedding module. The resulting representation allows for direct examination of motion primitive similarities, frequencies, and their distribution across different activities, providing unprecedented insights into model behavior. This comprehensive representation is subsequently processed by a Transformer backbone [48], enabling MoPFormer to effectively handle heterogeneous channel configurations and learn robust, interpretable features for diverse downstream tasks.
@@
  • We propose MoPFormer, an effective pre-training framework that achieves state-of-the-art performance.
@@
  • We extract motion primitives and provide a detailed interpretability analysis of them.
@@ 2 Related Work
Traditional human activity recognition approaches train model parameters on different datasets separately, including statistical feature extraction methods and deep learning approaches. Deep learning methods include CNN-based architectures such as CALANet [33], COA-HAR [50], MA-CNN [37], and SenseHAR [24], RNN-based approaches like DeepConvLSTM [30], and attention-based models such as THAT [26] and PA-HAR [51]. Recent research has proposed general-purpose time series models applicable to various tasks, including classification, such as TimesNet [49], TSLANet [17], and FITS [52]. However, these methods are typically trained and tested on splits from the original datasets and are limited in their ability to achieve cross-dataset generalization.
@@
Self-supervised human activity recognition methods perform representation learning through various approaches. Reconstruction-based methods such as TST [55] and TARNet [13] focus on rebuilding input signals. Contrastive learning methods such as TS2Vec [53], CL-HAR [35], DDLearn [36], TS-TCC [16], FOCAL [28], and ModCL [22] execute discriminative representation learning tasks. Other self-supervised objectives like BioBankSSL [55, 14] and Step2Heart [45] also contribute to learning time series representations. These methods typically learn feature representations on specific datasets and fine-tune classifier heads on the target dataset, but they can only perform on datasets with similar structures and still require labeled data from the target dataset.
@@
Research on sensor-based HAR for composite activity recognition is relatively limited [11]. For instance, Chen et al. decomposed composite activities into multiple simple activities [11], where each simple activity is represented by a sequence of sensor signal segments. These segments are first fed into a CNN to extract representations for identifying simple activities. Concurrently, the CNN-extracted features from all segments are passed to an LSTM network to achieve high-level semantic activity classification. Similarly, prior work [12] inferred composite activities using estimated activity sequences, where temporal correlations of simple activities were extracted for composite activity classification. Conversely, predicted composite activities were used to aid the derivation of simple activity sequences for the next time step, with predictions for both simple and composite activity sequences mutually updating during inference. While these studies present viable approaches for composite activity research, datasets with comprehensive, multi-level hierarchical labels for composite activities are scarce, and existing hierarchical annotations are often incomplete.
@@
The concept of motion primitives has gained traction in robotics research [41, 9]. These primitives are fundamental movement patterns that can be sequenced and combined to generate or understand a diverse repertoire of complex human activities. For instance, Calvo et al. utilized eight such primitive movements, including “static” and partial “squat”, to train Hidden Markov Models for activity classification. Similarly, Sanzari et al. developed a framework for automatically discovering human motion primitives from 3D pose data by optimizing “motion flux” and then clustering the results [41]. However, the application of motion primitives to enhance the interpretability of IMU-based HAR has seen limited progress. This limitation stems largely from the inherently non-intuitive nature of IMU sensor data, which makes it challenging to define and interpret primitives directly from these signals.
@@
Much of the research on interpretability in HAR has focused on attention mechanisms. The principle behind deep attention models is to weight input components, with the assumption that components assigned higher weights are more relevant to the recognition task and exert greater influence on the model’s decisions [43]. For example, Shen et al. [44] designed a segment-level attention method to determine which time periods contain more information; combined with a gated CNN, this segment-level attention can extract temporal dependencies. Zeng et al. [54] developed attention mechanisms from two perspectives: they first proposed sensor attention on the input to extract salient sensory patterns, and then applied temporal attention to LSTMs to filter out inactive data segments. Ma et al. [29] employed spatial and temporal attention mechanisms, extracting spatial dependencies by fusing patterns with self-attention. However, their interpretability often remains at a superficial level, indicating what the model focuses on, rather than why or how these features contribute to the recognition of complex, hierarchically structured activities.
@@ 3 Methodology
As illustrated in Fig. 2, the architecture of MoPFormer, comprises four modules: the motion primitive module, the context-aware embedding module, and two task heads. The motion primitive module first transforms multi-dimensional time-series data windows into index sequences from a fixed vocabulary. To further extract patterns, the context-aware embedding module extracts semantic information from contextual relationships and embeds high-dimensional vector sequences for the Transformer backbone. The Masked Auto-Encoding (MAE) head, as a pre-training task, strengthens the ability of the context-aware embedding module to understand motion primitives, while the Classification (CLS) head enables the Transformer backbone to learn effective features for classification.
Figure 2: Detailed illustration of key modules in our motion-centric framework. (a) Motion Primitive Module: raw data windows from multiple sensor channels are partitioned into motion primitives and processed through instance normalization to generate Vector Quantization (VQ) indices and statistical features. (b) Context-Aware Embedding Module: special tokens are inserted alongside masked motion embedding tokens ($M$) to form the complete input representation, which combines motion primitive embeddings, positional encodings, and statistical feature embeddings. (c) Task Heads: The diagram shows transformed features $X^{*}$ representing the corresponding position vectors after Transformer Encoder processing. The MAE head utilizes $M^{*}$ positions for masked token prediction during pretraining, while the trainable CLS head operates exclusively on the transformed [CLS] token representation for downstream task fine-tuning.
@@ 3.1 Motion Primitive Module
The Motion Primitive Module transforms raw data windows into discrete motion primitives, serving as the foundational, interpretable units for the motion-centric representation. For each data window of shape $T\times C$, where $T$ is the number of time steps and $C$ is the number of channels (e.g., accelerometer, gyroscope), we first partition the sequence into non-overlapping segments across each channel independently. Each channel is divided into segments of length $L$, resulting in $S=\lfloor T/L\rfloor$ segments per channel and a total of $S\times C$ segments per window.
Each segment undergoes a feature extraction process, where both low-level temporal patterns and statistical characteristics are computed. For each segment $s_{i}\in\mathbb{R}^{L}$ (a single-channel segment), we apply instance normalization in Eq. (1) to standardize the input as
@@
The normalized segments are then encoded into discrete motion primitives using a Vector Quantization (VQ) [47] approach. We maintain a learnable codebook $\mathcal{Z}=\{z_{1},z_{2},...,z_{K}\}$, which contains $K$ prototype vectors (or “motion primitives”), each of dimension $L$. As illustrated in Fig. 2(a), for each normalized segment $\hat{s}_{i}$, we compute its VQ index through Eq. (2) by finding the nearest prototype in the codebook:
@@ 4.1 Experimental Setup
We conducted extensive evaluations using six publicly available benchmark datasets: PAMAP2 [38, 39], DSADS [7, 1], MHealth [5, 6], Realworld [46], UCI-HAR [40, 2], and USC-HAD [56]. To ensure fair comparison with existing self-supervised methods, we pre-trained our model on five of these datasets while using the remaining dataset for evaluation, maintaining consistent experimental settings with other self-supervised approaches. All datasets are resampled to 100 Hz and segmented with a 500-sample window. To eliminate information leakage from overlapping windows, we adopt the setting by using strides equal to the window size for all labeled training segments. This strict non-overlapping approach may reduce absolute performance compared with more permissive settings, but ensures a more rigorous evaluation. We use accuracy and macro-F1 as evaluation metrics to assess model performance.
@@ 4.2 Results
BYOL [21]+perm_jit Acc 80.88 91.30 88.94 85.18 82.66 74.04 83.83
@@
BYOL [21]+Mixup Acc 82.38 93.87 90.06 88.80 85.31 75.59 86.00
@@
We compared MoPFormer with six different baseline approaches. Following the comprehensive evaluation in [35], we selected the BYOL framework [21], which demonstrated the best average performance among various contrastive learning frameworks. For time-domain and frequency-domain augmentations, we used perm_jit and lfc, respectively, which showed superior performance in their domains. As noted in [15], Mixup provides distinctive benefits compared to other temporal augmentation methods, so we included it as a baseline. ModCL [22] represents a state-of-the-art contrastive learning method that leverages both intra-modal and inter-modal consistency in wearable sensor data. We included it as a contemporary contrastive learning approach. For all contrastive learning methods, we maintained a consistent 0.2:0.8 train-test split ratio.
Additionally, we compared against recent supervised approaches, including TSLANet [17] (ICML 2024), a widely recognized foundational model in the time-series domain, and CALANet [34] (NeurIPS 2024), which was specifically designed for HAR tasks [8]. Detailed parameter configurations for all baseline models are provided in Appendix A.